Let Y_{1}, …, Y_{n} be independent and identically distributed random variables with finite μ = E(Y) and σ^{2} = V (Y). Let U_{n} = \frac{\overline{Y} −μ}{σ/\sqrt{n}}. Then as n → ∞, U_{n} converges to a standard normal random variable in distribution.